OpenStudy (anonymous):

Can you show me how to solve this? Solve 6.3 = 0.9y y =

OpenStudy (jhannybean):

divide both sides by 0.9

OpenStudy (anonymous):

6.3/0.9=7 0.9/0.9=1 is that the correct way?

OpenStudy (jhannybean):

yes but don't forget you still have your `y` on the right side.

OpenStudy (anonymous):

Okay I am not understanding how to find Y....sorry

OpenStudy (jhannybean):

Typo there.

OpenStudy (jhannybean):

\[\frac{6.3}{0.9} =\frac{0.9}{0.9}y\] For \(\frac{6.3}{0.9}\) you gt \(7\) For \(\frac{0.9}{0.9}\) you got \(1\)

OpenStudy (jhannybean):

\[7=(1)y\]

OpenStudy (jhannybean):

Understand?

OpenStudy (anonymous):

Yes now I do....thank you !!

OpenStudy (jhannybean):

No problem :)

OpenStudy (anonymous):

:)